# Discrete Approximate Circle Bundles

### v2

@@ A.1 The Group $\text{Homeo}(\mathbb{S}^{1})$
Suppose $\tilde{f}\in\widetilde{\text{Homeo}_{\pm}(\mathbb{S}^{1})}$ is $L$-Lipschitz. Then, $d_{\infty}(\tilde{f},T^{\pm}_{\tilde{f}(0)})\leq\frac{L-1}{L}$.
@@
Assume $\tilde{f}\in\widetilde{\text{Homeo}_{+}(\mathbb{S}^{1})}$ (the proof for $\tilde{f}\in\widetilde{\text{Homeo}_{-}(\mathbb{S}^{1})}$ is similar). For any $x\in(0,1)$, we have $\tilde{f}(x)-\tilde{f}(0)\leq Lx$ and $\tilde{f}(1)-\tilde{f}(x)\leq L(1-x)$. On the other hand, we also have $\tilde{f}(0)<\tilde{f}(x)<\tilde{f}(1)$ for any $x\in(0,1)$, so the point $(x,\tilde{f}(x))$ is constrained to lie in a polygonal region. In particular, the quantity $|y-(x+\tilde{f}(0))|$ is maximized at the corner points $(\frac{1}{L},\frac{L-1}{L})$ and $(1-\frac{1}{L},\frac{L-1}{L})$. It follows that $d_{\infty}(\tilde{f},T^{+}_{\tilde{f}(0)})\leq\frac{L}{L-1}$.
Proposition A.2.
[…]
@@
Assume $f$ is orientation-preserving (the proof when $f$ reverses orientation is similar), and let $\tilde{f}$ be any lift of $f$. Let $w,z\in\mathbb{S}^{1}$, and let $\tilde{w},\tilde{z}\in\mathbb{R}$ be lifts of $w$ and $z$ such that $d_{\mathbb{R}}(\tilde{w},\tilde{z})\leq\frac{1}{2}$ (such a pair of lifts always exists). We have
[…]
Proposition A.3.
Suppose $\tilde{f},\tilde{g}\in\widetilde{\text{Homeo}(\mathbb{S}^{1})}$ are lifts of $f,g\in\text{Homeo}(\mathbb{S}^{1})$, respectively. If $d_{\infty}(\tilde{f},\tilde{g})<\varepsilon$, then $d_{\infty}(f,g)<2\pi\varepsilon$.
Proof.
Choose any $z\in\mathbb{S}^{1}$ and let $\tilde{z}\in\mathbb{R}$ be any lift. Then,
$d_{\mathbb{S}^{1}}(f(z),g(z))\$$)\displaystyle=\ d_{\mathbb{S}^{1}}(\exp\circ\tilde{f}(\tilde{z}),\ \exp\circ\tilde{g}(\tilde{z}$))
$)\displaystyle\leq\ 2\pi d_{\mathbb{R}}(\tilde{f}(\tilde{z}),\ \tilde{g}(\tilde{z}$))
$<2\pi\varepsilon$
Proposition A.4.
If $f\in\text{Homeo}_{\pm}(\mathbb{S}^{1})$ is $L$-Lipschitz, then $d_{\infty}(f,R_{\log f(1)}^{\pm})\leq 2\pi(\frac{L-1}{L})$.
Proof.
By Proposition A.2, any lift $\tilde{f}$ of $f$ is $L$-Lipschitz, so by Proposition A.1, any lift satisfies $d_{\infty}(\tilde{f},T^{\pm}_{\tilde{f}(0)})<\frac{L-1}{L}$. Finally, using Proposition A.3, we obtain $d_{\infty}(f,R^{\pm}_{\log f(0)})=d_{\infty}(f,\exp\circ T_{\tilde{f}(0)})<2\pi(\frac{L-1}{L})$.
A.2 Circle Bundles Over $\mathbb{S}^{2}$ and $\mathbb{RP}^{2}$
In this section, we provide the details of the constructions and classifications of the two non-trivial circle bundle models described in the Implementations section. Since both models are associated with quotients of $SO(3)$, we begin by describing principal $\mathbb{S}^{1}$-bundle structures for $SO(3)$ and its universal cover $\mathbb{S}^{3}$.
@@ A.2 Circle Bundles Over $\mathbb{S}^{2}$ and $\mathbb{RP}^{2}$
Note, however, that the two principal circle bundles are not isomorphic. Indeed, Adj is not a homeomorphism, nor is it $\mathbb{S}^{1}$-equivariant (in fact, one has $\text{Adj}_{q\cdot e^{i\theta}}=\text{Adj}_{q}\cdot e^{i(2\theta)}$). The Hopf bundle famously has Euler number -1 with respect to the standard orientation on $\mathbb{S}^{2}$ (for a proof, see [Hus94]). The following proposition immediately implies that the Euler number of the principal bundle $(SO(3),\pi,\mathbb{S}^{2})$ is 2-2, and will also be useful for classifying the circle bundles from our models in Section 7:
@@
Let $\Omega\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{S}^{1}})$ be a cocycle representative of $\xi$ subordinate to some open cover $\mathcal{U}$ of $B$. By the results in Section 2.1, $\varphi_{p}^{*}\xi$ is a principal $\mathbb{S}^{1}$-bundle which is represented by $\varphi_{*}\Omega$. If $\theta\in\check{C}^{1}(\mathcal{U};\mathbb{R})$ is any lift of $\widetilde{\Omega}$ (i.e., $\Omega_{jk}(b)=e^{i\theta_{jk}(b)}$ for all $b\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$), then
$(\varphi_{*}\Omega)_{jk}(b)=\varphi\circ\Omega_{jk}(b)=\varphi(e^{i\theta_{jk}(b)})=e^{i(p\theta_{jk}(b))}$
for all $b\in U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$, so $p\theta$ is a lift of $\varphi_{*}\Omega$. The components of the Euler class $e(\varphi_{*}\xi)$ are then given by
$e(\varphi_{*}\xi)_{jkl}$ $=p\theta_{jk}(b)+p\theta_{kl}(b)-p\theta_{jl}(b)$
$=p(\theta_{jk}(b)+\theta_{kl}(b)-\theta_{jl}(b))$
$=pe(\xi)_{jkl}$
We conclude $e(\varphi_{*}\xi)=pe(\xi)$.
@@
Suppose $\xi=(E,\pi,B)$ is a principal $G$-bundle, and $E$ and $B$ have free, proper right actions by a group $H$ such that:
@@
Lastly, suppose that in addition to a $p$-fold rotational symmetry in the plane perpendicular to $e_{1}$, $x\in X$ also has an additional 2-fold rotational symmetry with respect to this plane (imagine a prism). Then $\text{stab}(x)$ is the non-abelian group generated by $\{R_{e_{1},\frac{2\pi}{p}},R_{e_{2},\pi}\}$, where $R_{e_{1},\theta}R_{e_{2},\pi}=R_{e_{2},\pi}R_{e_{1},-\theta}$. In particular, observe that $R_{e_{2},\pi}$ induces a free $\mathbb{Z}_{2}$-action on $SO(3)/\langle R_{e_{1},\frac{2\pi}{p}}\rangle$ given by $A]\cdot(-1)=[AR_{e_{2},\pi}]$. Moreover, one can easily verify that $\pi([A]\cdot(-1))=-\pi([A])$ and $([A]\cdot R_{e_{1},\theta})\cdot(-1)=([A]\cdot(-1))\cdot R_{e_{1},-\theta}$ for all $A\in SO(3)$. Thus, by Proposition A.7, the induced projection map $\pi^{\prime}:SO(3)/\langle R_{e_{1},\frac{2\pi}{p}},R_{e_{2},\pi}\rangle\to\mathbb{RP}^{2}$ given by $\pi^{\prime}([A])=[\pi(A)]$ is a non-orientable circle bundle. In fact, using the adjoint map to identify $\mathbb{S}^{3}/\mathbb{Z}_{2}$ with $SO(3)$, this bundle is canonically isomorphic to the circle bundle $\pi^{\prime}_{2p}:L(2p,1)/\mathbb{Z}_{2}\to\mathbb{RP}^{2}$ described in Corollary A.8 (note in particular that $\text{Adj}_{j}=R_{e_{2},\pi}$).

### v3

@@ A.1 Approximate Isometries Of The Circle
Suppose $h:\mathbb{S}^{1}\to\mathbb{S}^{1}$ is any function satisfying $\text{dist}(h)\leq\varepsilon$ for some $\varepsilon<\frac{\pi}{4}$. Then,
  1. 1.
    There exists some $\Omega\in O(2)$ such that $d_{\infty}(\Omega,h)\leq 2\varepsilon$,
  2. 2.
    The determinant of any such $\Omega$ is the same,
  3. 3.
    The map $d_{\infty}(h,\cdot):O(2)\to\mathbb{R}_{\geq 0}$ has a unique minimizer.
@@
First, observe that there is a unique $\Omega\in O(2)$ which satisfies $\Omega\circ h(1)=1$ and $d_{\mathbb{S}^{1}}(\Omega\circ h(i),i)\leq\varepsilon$. Let $\tilde{h}=\Omega\circ h$; to prove 1 above, it suffices to show $d_{\infty}(\tilde{h},\text{Id})\leq 2\varepsilon$. To see this, note that for any $z\in\mathbb{S}^{1}$, we have
$\left|d_{\mathbb{S}^{1}}(\tilde{h}(z),1)-d_{\mathbb{S}^{1}}(z,1)\right|\quad=\quad\left|d_{\mathbb{S}^{1}}(h(z),h(1))-d_{\mathbb{S}^{1}}(z,1)\right|\quad\leq\quad\varepsilon$ (39)
and
$\left|d_{\mathbb{S}^{1}}(\tilde{h}(z),i)-d_{\mathbb{S}^{1}}(z,i)\right|\quad$ $\leq\quad\left|d_{\mathbb{S}^{1}}(\tilde{h}(z),i)-d_{\mathbb{S}^{1}}(\tilde{h}(z),\tilde{h}(i))\right|\quad+\quad\left|d_{\mathbb{S}^{1}}(\tilde{h}(z),\tilde{h}(i))-d_{\mathbb{S}^{1}}(z,i)\right|$
$\leq\quad d_{\mathbb{S}^{1}}(i,\tilde{h}(i))\quad+\quad\left|d_{\mathbb{S}^{1}}(h(z),h(i))-d_{\mathbb{S}^{1}}(z,i)\right|$
$\leq\quad 2\varepsilon$
By Lemma A.4, $d_{\mathbb{S}^{1}}(\tilde{h}(z),z)\leq 2\varepsilon$ for all $z\in\mathbb{S}^{1}$, so $d_{\infty}(\tilde{h},\text{Id})\leq 2\varepsilon$.
2. Suppose $\Omega^{\prime}\in O(2)$ satisfies $d_{\infty}(h,\Omega^{\prime})\leq 2\varepsilon$ and $\det(\Omega^{\prime})\neq\det(\Omega)$. By the triangle inequality,
$d_{\infty}(\Omega^{\prime},\Omega)\quad\leq\quad d_{\infty}(\Omega^{\prime},h)\ +\ d_{\infty}(h,\Omega)\quad\leq\quad 4\varepsilon\quad<\quad\pi$
On the other hand, Lemma A.2 implies $d_{\infty}(\Omega^{\prime},\Omega^{-1})=\pi$, a contradiction. We conclude $\det(\Omega^{\prime})=\det(\Omega^{-1})=\det(\Omega)$.
3. Let $L_{h}:O(2)\to\mathbb{R}_{\geq 0}$ be defined by $L_{h}(\Omega)=d_{\infty}(h,\Omega)$. By compactness, at least one minimizer of $L_{h}$ exists. By 1, the set of minimizers is contained in $S_{\varepsilon}=L_{h}^{-1}([0,2\varepsilon])$, and 2 implies $S_{\varepsilon}$ is contained in a single connected component of $O(2)$. We assume minimizers of $L_{h}$ have positive determinant (if not, the same result is obtained by first pre-composing $h$ with any reflection in $O(2)$).
For $t\in[0,2\pi)$, define $s(t)\in[0,2\pi)$ such that $h(e^{it})=e^{is(t)}$, and let g(t)=s(t)tg(t)=s(t)-t. If $R_{\theta}$ minimizes $L_{h}$ for some $\theta\in[0,2\pi)$, then
$d_{\mathbb{S}^{1}}(e^{ig(t)},e^{i\theta})\quad=\quad d_{\mathbb{S}^{1}}(e^{i(s(t)-t)},e^{i\theta})\quad=\quad d_{\mathbb{S}^{1}}(e^{is(t)},e^{i(\theta+t)})\quad\leq\quad L_{h}(R_{\theta})\quad\leq\quad 4\varepsilon\quad<\quad\frac{\pi}{2}$
for all $t\in[0,2\pi)$. It follows that the image $e^{ig([0,2\pi))}$ is contained in an arc of diameter at most $4\varepsilon<\pi$ containing $e^{i\theta}$. The unique minimizer of $L_{h}$ is then $R_{\theta^{*}}$, where $e^{i\theta^{*}}$ is the (unique) midpoint of the smallest arc containing $e^{ig([0,2\pi))}$. From uniqueness of the minimizer, we conclude $\theta^{*}=\theta\mod 2\pi$.
We finally arrive at the main result of this section:
Lemma A.6.
Let $\pi:X\to B$ be a function between metric spaces. Suppose $\{\varphi_{j}:(\pi^{-1}(U_{j}),\pi,U_{j})\to(U_{j}\times\mathbb{S}^{1}_{r_{j}},p_{B},U_{j})\}_{j=1}^{2}$ are discrete $(r,\beta,\varepsilon)$-approximate local trivializations for some $r_{1},r_{2}\geq r$ and open sets $U_{1},U_{2}\subseteq B$ such that $\bar{B}_{\beta}(\pi(x))\subset U_{12}=U_{1}\cap U_{2}$ for some $x\in X$. Then,
$|r_{1}-r_{2}|\leq\frac{2}{\pi}(\varepsilon+\beta)$
and there exists some $\Omega_{12}\in O(2)$ such that
$|f_{1}(x)-\Omega_{12}f_{2}(x)|\quad\leq\quad\tilde{\tau}(\varphi_{1},\varphi_{2})$
for all $x\in\pi^{-1}(U_{12})$, where $f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}$ for $j=1,2$ and $\widetilde{\tau}(\varphi_{1},\varphi_{2})$ is defined by (14).
@@
For $j=1,2$, let $\psi_{j}:B\times\mathbb{S}^{1}\to X$ be an almost-inverse for $\varphi_{j}$. Let $\pi_{j}=p_{B}\circ\varphi_{j}$ and $f_{j}=p_{\mathbb{S}^{1}}\circ\varphi_{j}$. Choose some $x_{12}\in\pi^{-1}(U_{1}\cap U_{2})$ such that $\bar{B}_{\beta}(\pi(x_{12}))\subset U_{1}\cap U_{2}$, let $b_{12}=\pi(x_{12})$ and let $b_{j}=p_{B}\circ\varphi_{j}(x_{12})$ for $j=1,2$. For any $z\in\mathbb{S}^{1}$, we have $d_{B}(\pi\circ\psi_{2}(b_{2},z),b_{12})\leq\frac{\beta}{2}$, so $\psi_{2}(b_{2},z)\in\bar{B}_{\frac{\beta}{2}}(b_{12})\subseteq\bar{B}_{\beta}(b_{12})\subseteq U_{jk}$. We therefore have a well-defined function $\Lambda_{12}:\mathbb{S}^{1}_{r_{2}}\to\mathbb{S}^{1}_{r_{1}}$ given by $\Lambda_{12}(z)=f_{1}\circ\psi_{2}(b_{2},z)$ for all $z\in\mathbb{S}^{1}$.
We now estimate $\text{dist}(\Lambda_{12})$ in order to invoke Proposition A.5. Let $T_{12}:\mathbb{S}^{1}\to U_{1}\times\mathbb{S}^{1}$ be defined by $T_{12}(z)=\varphi_{1}\circ\psi_{2}(b_{2},z)$ for all $z\in\mathbb{S}^{1}$. Note that for any $z\in\mathbb{S}^{1}$,
$d_{B}(\pi\circ\psi_{2}(b_{2},z),b_{12})\quad\leq\quad d_{B}(\pi\circ\psi_{2}(b_{2},z),b_{2})+d_{B}(b_{2},b_{12})\quad\leq\quad\beta$
so $\psi_{2}(b_{2},z)\in\bar{B}_{2\beta}(b_{12})\subset\pi^{-1}(U_{1})$ and $T_{12}(z)$ is well-defined. Moreover, for any $z,z^{\prime}\in\mathbb{S}^{1}$,
$d_{B}(p_{B}\circ T_{12}(z),\ p_{B}\circ T_{12}(z^{\prime}))\$$=\ d_{B}(\pi_{1}\circ\psi_{2}(b_{2},z),\ \pi_{1}\circ\psi_{2}(b_{2},z^{\prime}))$
$\leq d_{B}(\pi_{1}\circ\psi_{2}(b_{2},z),\ \pi\circ\psi_{2}(b_{2},z))$
$+d_{B}(\pi\circ\psi_{2}(b_{2},z),b_{2})$
$+d_{B}(b_{2},\pi\circ\psi_{2}(b_{2},z^{\prime}))$
$+d_{B}(\pi\circ\psi_{2}(b_{2},z^{\prime}),\pi_{1}\circ\psi_{2}(b_{2},z^{\prime}))$
$\leq 2\beta$
Since $\text{dist}(\varphi_{1}),\text{dist}(\psi_{2})\leq\varepsilon$, we have
$|d_{\infty}^{r_{1}}(T_{12}(z),T_{12}(z^{\prime}))-d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})|\quad\leq\quad 2\varepsilon$
which in particular implies
$d_{\mathbb{S}^{1}}^{r_{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))\quad\leq\quad d^{r_{1}}_{\infty}(T_{12}(z),T_{12}(z^{\prime}))\quad\leq\quad d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})+2\varepsilon$ (40)
for any $z,z^{\prime}\in\mathbb{S}^{1}$. On the other hand, since
$d^{r_{1}}_{\infty}(T_{12}(z),T_{12}(z^{\prime}))\quad=\quad\max\{d_{B}(p_{B}\circ T_{12}(z),\ p_{B}\circ T_{12}(z^{\prime})),\hskip 14.22636ptd^{r_{1}}_{\mathbb{S}^{1}}(\Lambda_{12}(z),\ \Lambda_{12}(z^{\prime}))\}$
we can apply the algebraic identity $v\geq\max\{u,v\}-u$ (for any $u,v\geq 0$) to obtain
$d_{\mathbb{S}^{1}}^{r_{1}}(\Lambda_{12}(z),\ \Lambda_{12}(z^{\prime}))\$$\geq\ d_{\infty}^{r_{1}}(T_{12}(z),T_{12}(z^{\prime}))-d_{B}(p_{B}\circ T_{12}(z),\ p_{B}\circ T_{12}(z^{\prime}))$
$\geq(d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})-2\varepsilon)-2\beta$
$=d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})-2(\varepsilon+\beta)$
Combining with (40), we conclude $\text{dist}(\Lambda_{12})\leq 2(\varepsilon+\beta)$ as a function from $\mathbb{S}^{1}_{r_{2}}$ to $\mathbb{S}^{1}_{r_{1}}$. As an immediate consequence, we find that for any $z,z^{\prime}\in\mathbb{S}^{1}$,
$\pi r_{2}-2(\varepsilon+\beta)\quad\leq\quad d_{\mathbb{S}^{1}}^{r_{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))\quad\leq\quad\pi r_{1}$
which implies $\pi r_{2}-2(\varepsilon+\beta)\leq\pi r_{1}$, or $\pi(r_{2}-r_{1})\leq 2(\varepsilon+\beta)$. Swapping the roles of $r_{1}$ and $r_{2}$ yields $\pi(r_{1}-r_{2})\leq 2(\varepsilon+\beta)$, so
$|r_{2}-r_{1}|\quad\leq\quad\frac{2}{\pi}(\varepsilon+\beta)$
Now, observe
$2(\varepsilon+\beta)\quad\geq\quad\left|d_{\mathbb{S}^{1}}^{r_{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))-d_{\mathbb{S}^{1}}^{r_{2}}(z,z^{\prime})\right|\quad=\quad\left|r_{1}d_{\mathbb{S}^{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))-r_{2}d_{\mathbb{S}^{1}}(z,z^{\prime})\right|$
which implies
$\left|d_{\mathbb{S}^{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))-d_{\mathbb{S}^{1}}(z,z^{\prime})\right|\quad$ $\leq\quad\left|d_{\mathbb{S}^{1}}(\Lambda_{12}(z),\Lambda_{12}(z^{\prime}))-\frac{r_{2}}{r_{1}}\ d_{\mathbb{S}^{1}}(z,z^{\prime})\right|+\left|\frac{r_{2}}{r_{1}}d_{\mathbb{S}^{1}}(z,z^{\prime})-d_{\mathbb{S}^{1}}(z,z^{\prime})\right|$
$\leq\quad\frac{2}{r_{1}}(\varepsilon+\beta)+\frac{d_{\mathbb{S}^{1}}(z,z^{\prime})}{r_{1}}|r_{2}-r_{1}|$
$\leq\quad\frac{2}{r_{1}}(\varepsilon+\beta)+\frac{\pi}{r_{1}}|r_{2}-r_{1}|$
$\leq\quad\frac{4}{r_{1}}(\varepsilon+\beta)$
Since $r_{1}>\frac{16}{\pi}(\varepsilon+\beta)$ by definition of a discrete $(r,\beta,\varepsilon)$-approximate local trivialization, the inequality above implies that, as a function from $\mathbb{S}^{1}$ to $\mathbb{S}^{1}$ (with the usual unscaled geodesic distance), $\text{dist}(\Lambda_{12})<\frac{\pi}{4}$. Proposition A.5 then implies that the loss function $\Omega\longmapsto d_{\infty}(\Lambda_{12},\Omega)$ has a unique minimizer $\Omega_{12}$ in $O(2)$, and that $d_{\infty}(\Lambda_{12},\Omega_{12})\leq\frac{8}{r_{1}}(\varepsilon+\beta)$.
Now, let $\tilde{x}\in\pi^{-1}(U_{12})$ be arbitrary. Let $\varphi_{2}(\tilde{x})=(\tilde{b},\tilde{z})\in U_{2}\times\mathbb{S}^{1}_{r_{2}}$ and let $\tilde{y}=\psi_{2}(b_{2},\tilde{z}_{2})\in\pi^{-1}(U_{12})$. Observe,
$d_{B}(\tilde{b}_{2},b_{2})\quad$ $\leq\quad d_{B}(\tilde{b}_{2},\pi(\tilde{x}))+d_{B}(\pi(\tilde{x}),\pi(x_{12}))+d_{B}(\pi(x_{12}),b_{2})$
$\leq\quad\frac{\beta}{2}+\text{diam}(U_{12})+\frac{\beta}{2}$
$=\quad\text{diam}(U_{12})+\beta$
By the codistortion property of $\varphi_{2}$ and $\psi_{2}$,
$d_{X}(\tilde{x},\tilde{y})\quad\leq\quad d_{\infty}^{r_{2}}(\varphi_{2}(\tilde{x}),(b_{2},\tilde{z}_{2}))+\varepsilon\quad\leq\quad d_{B}(\tilde{b}_{2},b_{2})+\varepsilon\quad\leq\quad\text{diam}(U_{12})+\varepsilon+\beta$
and since $\text{dist}(\varphi_{1})\leq\varepsilon$,
$d_{\mathbb{S}^{1}}^{r_{1}}(f_{1}(\tilde{x}),f_{1}(\tilde{y}))\quad\leq\quad d_{\infty}^{r_{1}}(\varphi_{1}(\tilde{x}),\varphi_{1}(\tilde{y}))\quad\leq\quad d_{X}(\tilde{x},\tilde{y})+\varepsilon\quad\leq\quad\text{diam}(U_{12})+2\varepsilon+\beta$
or, equivalently,
$d_{\mathbb{S}^{1}}(f_{1}(\tilde{x}),f_{1}(\tilde{y}))\quad\leq\quad\frac{1}{r_{1}}\left(\text{diam}(U_{12})+2\varepsilon+\beta\right)$
Finally, noting that $)f_{1}(\tilde{y})=\Lambda_{12}(f_{2}(\tilde{x}$)), we obtain
$)\displaystyle d_{\mathbb{S}^{1}}(f_{1}(\tilde{x}),\Omega_{12}f_{2}(\tilde{x}$$)\quad$ $)\displaystyle\leq\quad d_{\mathbb{S}^{1}}(f_{1}(\tilde{x}),f_{1}(\tilde{y}))+d_{\mathbb{S}^{1}}(f_{1}(\tilde{y}),\Omega_{12}f_{2}(\tilde{x}$))
$)\displaystyle\leq\quad\frac{1}{r_{1}}(\text{diam}(U_{12})+2\varepsilon+\beta)+d_{\mathbb{S}^{1}}(\Lambda_{12}(f_{2}(\tilde{x})),\Omega_{12}f_{2}(\tilde{x}$))
$\leq\quad\frac{1}{r_{1}}(\text{diam}(U_{12})+2\varepsilon+\beta+8(\varepsilon+\beta))$
$=\quad\frac{1}{r_{1}}(\text{diam}(U_{12})+10\varepsilon+9\beta)$
$\leq\quad\tau(\varphi_{1},\varphi_{2})$
A.2 Circle Bundles Over $\mathbb{RP}^{2}$
In this section, we provide some details of the construction and classification of the circle bundle underlying the dataset used in Section 7.3. The bundle is obtained as a quotient of the classical principal $\mathbb{S}^{1}$-bundle over $\mathbb{S}^{2}$ with total space $SO(3)$, briefly described below:
@@ A.2 Circle Bundles Over $\mathbb{RP}^{2}$
Note, however, that the two principal circle bundles are not isomorphic. Indeed, Adj is not a homeomorphism, nor is it $\mathbb{S}^{1}$-equivariant (in fact, one has $\text{Adj}_{q\cdot e^{i\theta}}=\text{Adj}_{q}\cdot e^{i(2\theta)}$). The Hopf bundle famously has Euler number -1 with respect to the standard orientation on $\mathbb{S}^{2}$ (for a proof, see [12]). The following proposition immediately implies that the Euler number of the principal bundle $(SO(3),\pi,\mathbb{S}^{2})$ is 2-2, and will also be useful for classifying the circle bundle model underlying the density dataset in Section 7:
@@
See, for example, Chapter 4 of  [12].
@@
Suppose $\xi=(E,\pi,B)$ is a principal $G$-bundle, and $E$ and $B$ have free, proper right actions by a topological group $H$ such that:
@@
Lastly, suppose that in addition to a $p$-fold rotational symmetry in the plane perpendicular to $e_{1}$, $x\in X$ also has an additional 2-fold rotational symmetry with respect to this plane (as for the triangle prism density functions in Section 7.3). Then $\text{stab}(x)$ is the non-abelian group generated by $\{R_{e_{1},\frac{2\pi}{p}},R_{e_{2},\pi}\}$, where $R_{e_{1},\theta}R_{e_{2},\pi}=R_{e_{2},\pi}R_{e_{1},-\theta}$. In particular, observe that $R_{e_{2},\pi}$ induces a free $\mathbb{Z}_{2}$-action on $SO(3)/\langle R_{e_{1},\frac{2\pi}{p}}\rangle$ given by $A]\cdot(-1)=[AR_{e_{2},\pi}]$. Moreover, one can easily verify that $\pi([A]\cdot(-1))=-\pi([A])$ and $([A]\cdot R_{e_{1},\theta})\cdot(-1)=([A]\cdot(-1))\cdot R_{e_{1},-\theta}$ for all $A\in SO(3)$. Thus, by Proposition A.9, the induced projection map $\pi^{\prime}:SO(3)/\langle R_{e_{1},\frac{2\pi}{p}},R_{e_{2},\pi}\rangle\to\mathbb{RP}^{2}$ given by $\pi^{\prime}([A])=[\pi(A)]$ is a non-orientable circle bundle. In fact, using the adjoint map to identify $\mathbb{S}^{3}/\mathbb{Z}_{2}$ with $SO(3)$, this bundle is canonically isomorphic to the circle bundle $\pi^{\prime}_{2p}:L(2p,1)/\mathbb{Z}_{2}\to\mathbb{RP}^{2}$ described in Corollary A.10 (note in particular that $\text{Adj}_{j}=R_{e_{2},\pi}$).